8) A social psychologist studying sources of knowledge randomly assigned 82 volunteers to one of two experimental groups. 61 were instructed to get their news for a month only from television, and 21 were instructed to get their news for a month only from the Internet. After the month was up, all participants were tested on their knowledge of several political issues. The researcher did not have a prediction as to which news source would make people more knowledgeable. These were the results of the study:

TV group: M1=24, SS1=240

Internet group: M2=26, SS2=120

a) What is the null and alternative hypothesis?

b)Compute tobt?

c)Compute df?

d)Given the α=0.01, compute the negative value of tcrit?

e)Is tobt significant?
Answer #1

Ans:

a)

Ho : μι Σμ2

HQ: M1 <H2

b)

pooled standard deviation=SQRT((240+120)/(61+21-2))=2.121

standard error for difference=2.121*sqrt((1/61)+(1/21))=0.5367

Test statistic:

tobt=(24-26)/0.5367

tobt=-3.73

c)df=61+21-2=80

d)critical t value(left tailed)=-2.374

e)As,tobt is less than critical t value,So,

yes,tobt is significant.